OpenStudy (anonymous):

What is the product of 5the square root of 25 times d. • 2dthe square root of 3.? Simplify if possible. 5dthe square root of 3. 10dthe square root of 5 times d. 25dthe square root of 6 times d. 50dthe square root of 3 times d.

OpenStudy (anonymous):

problem here ' 2dthe '

OpenStudy (anonymous):

it's D I think

OpenStudy (anonymous):

\[5\sqrt{25}= 5\times5 =25 d \times 2\sqrt{3} = 50\sqrt{3}d\]

OpenStudy (anonymous):

\[sorry it's 2\sqrt{3d} \times \sqrt{d} 25 = 50 d \sqrt{3}\]

OpenStudy (anonymous):

thank you!